Consider a harmonic oscillator with Hamiltonian given by ?=(p^2/2m)+(1/2)X^2 = (a+)(a-)+(1/2)

The current system state is the superposition of the lowest and next-to-lowest energy eigenstates that gives the most negative possible value for the average position, use raising and lowering operators to derive the average momentum for this state.

then, simplify using ħ = ? = 1
Answer #1

H = 2 P 2m + p=ijn 2 1 2 = (ata + 1) <p> = (n/Pln+1) nhw (ata) c nhw (nlat-a/n+1) - domku { <nla lnt) - <n/aIn+1)] in Int2/nplz like